# O0.45cmO0.1cm =

### v1

@@ 4 Effective Borel Conjecture
    $X\subseteq\{R(\delta_{k})|k\in\mathbb{N}\}$.
@@
In the original definition it was required that $X=\{R(\delta_{k})|k\in\mathbb{N}\}$. For our purposes, we want all subsets of countable sets to be countable. Otherwise it is easy to come up with a set of languages $L_{i}$ with $|L_{i}|=1$ whose union is not $\mathrm{\Delta}$-countable while having $\mathrm{\Delta}$-strong measure zero.
@@
If $X$ is $\Delta$-countable then $X\in\mathrm{\Delta}\operatorname{-SMZ}$,
@@
To see that $d$ is $\mathrm{\Delta}_{RCA}$-computable let $d^{\prime}:\mathbb{N}\times\{0,1\}^{*}\to[0,\infty)\in\mathrm{\Delta}_{RCA}$ be the function
@@
While it is known that strong measure zero implying countability is independent of ZFC in the classical ($\mathrm{\Delta}=\mathrm{all}$) case, we will see the answers in the effective versions depend on the level of effectivity.
@@ 4.1 Time and Space Bounded Strong Measure Zero
Let $f(n)$ be a function and $S\in\mathbf{C}$ be a sequence with $S\not=R(\delta)$ for any $\delta\in\mathrm{DTIME}(f(n))$. Then there is an odds function $O$ such that no odds functional $F\in\mathrm{DTIME}(f(n))$ succeeds on $S$. Similarly for DSPACE.
@@
To see that this is possible at each stage, note that the oracle $O_{s-1}$ contains a finite amount of information. Moreover, as $F_{i}$ is resource bounded it has to eventually produce output even though $O_{s-1}$ is eventually all $1$’s and not acceptable. For any $i\leq s$ there must be infinitely many $k\in\mathbb{N}$ where $F_{i}^{O_{s-1}}(S[0\dots k+1])\leq\frac{1}{2}F_{i}^{O_{s-1}}(S[0\dots k])$ as otherwise a function could be created in $\mathrm{DTIME}(f(n))$ that computes $S$. There are also only $s-1$ possible naturals $k$ such that $F_{i}^{O_{s-1}}(S[0\dots k+1])>F_{i}^{O_{s-1}}(S[0\dots k])$. Hence ${F_{i}^{O_{s-1}}(S[0\dots n])\leq\frac{1}{2}}$ for all sufficiently large $n$.
@@
    For any $S\in R(\mathrm{\Delta})$, there is an $i$ such that $S\in X_{i}$.
@@
Now for each $f_{i}$ create a function $\delta_{i}:\mathbb{N}\times\{0,1\}^{*}\to\{0,1\}^{*}$ such that ${\delta_{i,k}(w)=\delta_{i}(k,w)}$ does the following. Let $k=\langle j,c\rangle$ for some $j,c\in\mathbb{N}$. Then run the Turing machine $M_{j}$ on input $s_{|w|}$ for $f_{i}(|w|)+c$ steps. If it halts and accepts then output $w1$, otherwise $w0$. Then each $\delta_{i,k}$ is a constructor and $\{R(\delta)|\delta\in\mathrm{DTIME}(f_{i})\}=\{R(\delta_{i,k})|k\in\mathbb{N}\}$. Hence, letting $X_{i}=\{R(\delta_{i,k})|k\in\mathbb{N}\}$ the lemma holds. ∎
@@
We prove the contrapositive, suppose $X$ is not $\mathrm{\Delta}$-countable and let $F$ be any odds functional in $\mathrm{\Delta}$. Let $f_{0},f_{1},f_{2},\dots$ be a sequence of functions that define $\mathrm{\Delta}$. Then $F\in\mathrm{DTIME}(f_{m})$ on odds oracles $O$ that have range $\{1,2\}$ for some $m\in\mathbb{N}$. By Lemma 4.4, there must be some $S\in X$ such that $S\neq R(\delta)$ for any $\delta\in\mathrm{DTIME}(f_{m})$. Thus, by Lemma 4.3 there is an $O$ which $F$ does not succeed on. ∎
@@
The hypothesis of NP not having measure zero in E and the weaker hypothesis of NP not having dimension zero in E have lead to many interesting consequences. For specific definitions and examples see [19, 23]. For the case of strong measure zero, we get an even weaker hypothesis that gives a tight bound on the complexity of NP unlike in the other two cases.
@@ 4.2 Computable Strong Measure Zero
$f(n)=\max\{t\in\mathbb{N}|\exists k\in\mathbb{N},k\leq n\text{ and }M_{k}(k)\text{ halts in exactly }t\text{ steps}\}.$
Then the language $A=\mathrm{Graph}(f)=\{\langle n,f(n)\rangle|n\in\mathbb{N}\}$ is almost constructible, but not decidable.
@@
Let $S$ be as given with $\delta$ being a witness to $S$ being almost constructible. Define a computable predictor functional that works as follows on prediction order $h$. By Corollary 3.8 we can assume $h$ is strictly increasing. The following recursive algorithm initiated with $w=\lambda$ and $L=0$ will give the values of a superpredictor $\pi$.
@@
7:  Recursively start Predictor on inputs ($wa,L+\log\frac{1}{p}$) ($wx,L+\log\frac{1}{1-p}$)
@@
Since the singleton set $\{S\}$ is not computably countable for any undecidable $S$, we have the following.
@@ 5 Algorithmic Strong Measure Zero
There is a universal lower semicomputable continuous semimeasure $\mathbf{M}$. That is, for any lower semicomputable continuous semimeasure $\mu$ there is a constant $c>0$ such that
@@
is a universal lower semicomputable supermartingale. That is, for any lower semicomputable supermartingale f there is a constant $c>0$ with
@@
Both of these theorems relativize to any oracle giving us the following
@@
There is an universal algorithmic functional $\mathbf{F}$ defined by ${\mathbf{F}(h)=\mathbf{d}^{h}}$ for all gauge functions $h:\mathbb{N}\to\mathbb{Q}^{+}$. Specifically, a set $X\subseteq\mathbf{C}$ has algorithmic strong measure zero if and only if for every order function $h$ and every $S\in X$,
$\limsup_{n\to\infty}\frac{\mathbf{d}^{h}(S[0\dots n])}{2^{n}h(n)}=\infty.$
@@
$\mathrm{KM}^{h}(x)=\log\frac{1}{\mathbf{M^{h}(x)}}$
@@
A sequence $S$ has algorithmic strong measure zero if and only if for every gauge function $h$ there are infinitely many $n$ such that
$\mathrm{KM}^{h}(S[0\dots n])\leq\log\frac{1}{h(n)}.$ (5.1)
@@
Let $S$ be any sequence and $h$ be a gauge function. We have that $S$ has algorithmic strong measure zero if and only if
$\limsup_{n\to\infty}\frac{\mathbf{d}^{h}(S[0\dots n])}{2^{n}h(n)}=\infty$
$\iff$ $\limsup_{n\to\infty}\frac{\mathbf{M}^{h}(S[0\dots n])}{h(n)}=\infty$
$\iff$ $\limsup_{n\to\infty}\frac{2^{-\mathrm{KM}^{h}(S[0\dots n])}}{h(n)}=\infty.$ (5.2)
[…]
$\mathrm{KM}^{h}(S[0\dots n])\leq\log\frac{1}{h(n)^{2}}+c.$
@@
$\frac{2^{-\mathrm{KM}^{h}(S[0\dots n])}}{h(n)^{2}}\geq 2^{-c}.$
[…]
For the other direction we will prove the contrapositive. Suppose there is gauge $h$ such that $\mathrm{KM}^{h}(S[0\dots n])>\log\frac{1}{h(n)}$ for all sufficiently large $n$. Then we have
$\frac{2^{-\mathrm{KM}^{h}(S[0\dots n])}}{h(n)}\leq 1$
[…]
@@
  • $X\in\mathbf{C}$ is in NCR if it its not $\mu$-random for any continuous probability measure $\mu$.
@@
First assume $S$ has algorithmic strong measure zero and consider any $Z$. Then for any unbounded increasing $h$ computable from $Z$ we have that ${g(n)=2^{-h(n)}}$ is a gauge also computable from $Z$. As $S$ has algorithmic strong measure zero, by Lemma 5.4 there is a constant $c$ and infinitely many n with
@@
Note that if there were only a finitely many $n$ then an $h^{\prime}$ could be computed by changing finitely many values so that S is strongly $h^{\prime}$-complex. Thus, there are infinitely many $n$ with
@@ 6 Effective coverings and a Correspondence Principle
Let $\mathrm{\Delta}\in\mathrm{\Gamma}_{RCA}$, $A\subseteq\mathbb{N}$ and $X\subseteq\mathbf{C}$. A function ${f^{A}:\mathbb{N}\to\{0,1\}^{*}}$ is a $\mathrm{\Delta}$-$A$ covering of $X$ if $f\in\mathrm{\Delta}$ with oracle $A:\mathbb{N}\to\{0,1\}$ satisfies the following.
@@
We refer to $\{f^{A}(n)|n\in A\}$ of such an $f$ as a $\mathrm{\Delta}$-$A$ cover of $X$ and say that $X$ is strongly $\mathrm{\Delta}$-coverable if there is a functional $F\in\mathrm{\Delta}$ such that $F(A)$ is a $\mathrm{\Delta}$-$A$ covering of $X$ for all infinite $A\subseteq\mathbb{N}$.
@@
4:  Run $F^{O}$ and for each $w$ with $|w|=f_{A}(n)$ where $F^{O}(w)\geq 1$ is found set $c=c-1$.
@@
[…]
@@
$f_{O}(n)=\min\{m\in\mathbb{N}|\prod_{i=0}^{m-1}O(w[0\dots i])\geq n\text{ for all }w\in\{0,1\}^{m}\}$
@@
Then $d_{i}$ is a $\Delta$-computable $O$-martingale. We have
@@
Let $\mathrm{\Delta}\in\mathrm{\Gamma}$. Then for any $X\subseteq\mathbf{C}$ there is an odds functional $F$ succeeding on $X$ if and only if there is an odds functional $F^{\prime}$ succeeding on $X$ where $F^{\prime}(O)$ is an odds martingale for all odds functions $O$. Moreover, there is a universal algorithmic martingale functional.
@@
  • $X\in\mathrm{\Sigma}_{2}^{0}$ if there is a computable function $f:\mathbb{N}\times\{0,1\}^{*}\to\{0,1\}$ where ${X=\bigcup_{i\in\mathbb{N}}X_{i}}$ for $X_{i}$ being the $\mathrm{\Pi}_{1}^{0}$ set defined by the tree ${T_{i}=\{w|f(i,w)=1\}}$.
@@
Moreover 1 and 2 are equivalent for any union of $\mathrm{\Pi}_{1}^{0}$ sets.
@@
For any computable ordinal $\alpha$ there is a $x$ with $x\equiv_{\mathrm{T}}\emptyset^{(\alpha)}$ and $\{x\}$ having computable strong measure zero.

### v2

@@ 4 Effective Borel Conjecture
    $X\subseteq\{R(\delta_{k})\mid k\in\mathbb{N}\}$.
@@
In the original definition it was required that $X=\{R(\delta_{k})\mid k\in\mathbb{N}\}$. For our purposes, we want all subsets of countable sets to be countable. Otherwise it is easy to come up with a set of languages $L_{i}$ with $|L_{i}|=1$ whose union is not $\mathrm{\Delta}$-countable while having $\mathrm{\Delta}$-strong measure zero.
@@
Let $\mathrm{\Delta}\in\mathrm{\Gamma}$. If $X$ is $\Delta$-countable then $X\in\mathrm{\Delta}\operatorname{-SMZ}$.
@@
To see that $d$ is $\mathrm{\Delta}$-computable let $d^{\prime}:\mathbb{N}\times\{0,1\}^{*}\to[0,\infty)\in\mathrm{\Delta}$ be the function
@@
While it is known that strong measure zero implying countability is independent of ZFC in the classical case, we will see the answers in the effective versions depend on the level of effectivity.
@@ 4.1 Time and Space Bounded Strong Measure Zero
Let $f(n)$ be a function and $S\in\mathbf{C}$ be a sequence with $S\not=R(\delta)$ for every
$\delta\in\mathrm{DTIME}(f(n))$}. Then there is an odds function $O$ such that no odds functional
$F\in\mathrm{DTIME}(f(n))$ succeeds on $S$. Similarly for DSPACE.
@@
To see that this is possible at each stage, note that the oracle $O_{s-1}$ contains a finite amount of information. Moreover, as $F_{i}$ is resource bounded it has to eventually produce output even though $O_{s-1}$ is eventually all $1$’s and not acceptable. for every $i\leq s$ there must be infinitely many $k\in\mathbb{N}$ where $F_{i}^{O_{s-1}}(S[0\dots k+1])\leq\frac{1}{2}F_{i}^{O_{s-1}}(S[0\dots k])$ as otherwise a function could be created in $\mathrm{DTIME}(f(n))$ that computes $S$. There are also only s1s-1 possible naturals $k$ such that $F_{i}^{O_{s-1}}(S[0\dots k+1])>F_{i}^{O_{s-1}}(S[0\dots k])$. Hence $F_{i}^{O_{s-1}}(S[0\dots n])\leq\frac{1}{2}$} for all sufficiently large $n$.
@@
    for every $S\in R(\mathrm{\Delta})$, there is an $i$ such that $S\in X_{i}$.
@@
Now for each $f_{i}$ create a function $\delta_{i}:\mathbb{N}\times\{0,1\}^{*}\to\{0,1\}^{*}$ such that $\delta_{i,k}(w)=\delta_{i}(k,w)$} does the following. Let $k=\langle j,c\rangle$ for some $j,c\in\mathbb{N}$. Then run the Turing machine $M_{j}$ on input $s_{|w|}$ for $f_{i}(|w|)+c$ steps. If it halts and accepts then output $w1$, otherwise $w0$. Then each $\delta_{i,k}$ is a constructor and $\{R(\delta)\mid\delta\in\mathrm{DTIME}(f_{i})\}=\{R(\delta_{i,k})\mid k\in\mathbb{N}\}$. Hence, letting $X_{i}=\{R(\delta_{i,k})\mid k\in\mathbb{N}\}$ the lemma holds. ∎
@@
We prove the contrapositive, suppose $X$ is not $\mathrm{\Delta}$-countable and Let $f_{0},f_{1},f_{2},\dots$ be a sequence of functions that define $\mathrm{\Delta}$. By Lemma 4.4, for every $m\in\mathbb{N}$ there must be some $S\in X$ such that $S\neq R(\delta)$ for every $\delta\in\mathrm{DTIME}(f_{m})$. Thus, by Lemma 4.3 there is an $O$ that no odds functional in $\mathrm{DTIME}(f_{m})$ succeeds on for $S$. ∎
@@
The hypothesis of NP not having measure zero in E and the weaker hypothesis of NP not having dimension zero in E have led to many interesting consequences. For specific definitions and examples see [23, 19]. For the case of strong measure zero, we get an even weaker hypothesis that gives a tight bound on the complexity of NP unlike in the other two cases.
@@ 4.2 Computable Strong Measure Zero
$f(n)=\max\{t\in\mathbb{N}\mid\exists k\in\mathbb{N},k\leq n\text{ and }M_{k}(k)\text{ halts in exactly }t\text{ steps}\}.$
Then the language $A=\mathrm{Graph}(f)=\{\langle n,f(n)\rangle\mid n\in\mathbb{N}\}$ is almost constructible, but not decidable.
@@
Let $S$ be as given with $\delta$ being a witness to $S$ being almost constructible. Define a computable predictor functional that works as follows on prediction order $h$. By Observation 3.8 we can assume $h$ is strictly increasing. The following recursive algorithm initiated with $w=\lambda$ and $L=0$ will give the values of a superpredictor $\pi$.
@@
7: Recursively start Predictor on inputs ($wb,L+\log\frac{1}{p}$) ($wx,L+\log\frac{1}{1-p}$)
@@
Since the singleton set $\{S\}$ is not computably countable for every undecidable $S$, we have the following.
@@ 5 Algorithmic Strong Measure Zero
There is a universal lower semicomputable continuous semimeasure $\mathbf{M}$. That is, for every lower semicomputable continuous semimeasure $\mu$ there is a constant $c>0$ such that
@@
is a universal lower semicomputable supermartingale. That is, for every lower semicomputable supermartingale f there is a constant $c>0$ with
@@
Both of these theorems relativize to any oracle giving us the following.
@@
There is an universal algorithmic functional $\mathbf{F}$ defined by $\mathbf{F}(g)=\mathbf{d}^{g}$} for all gauge functions $g:\mathbb{N}\to\mathbb{Q}^{+}$. Specifically, a set $X\subseteq\mathbf{C}$ has algorithmic strong measure zero if and only if for every gauge function $g$ and every $S\in X$,
$\limsup_{n\to\infty}\frac{\mathbf{d}^{g}(S[0\dots n])}{2^{n}g(n)}=\infty.$
@@
$\mathrm{KM}^{g}(x)=\log\frac{1}{\mathbf{M}^{g}(x)}$
@@
A sequence $S$ has algorithmic strong measure zero if and only if for every gauge function $g$ there are infinitely many $n$ such that
$\mathrm{KM}^{g}(S[0\dots n])\leq\log\frac{1}{g(n)}.$ (5.1)
@@
Let $S$ be any sequence and $g$ be a gauge function. We have that $S$ has algorithmic strong measure zero if and only if
$\limsup_{n\to\infty}\frac{\mathbf{d}^{g}(S[0\dots n])}{2^{n}g(n)}=\infty$
$\iff$ $\limsup_{n\to\infty}\frac{\mathbf{M}^{g}(S[0\dots n])}{g(n)}=\infty$
$\iff$ $\limsup_{n\to\infty}\frac{2^{-\mathrm{KM}^{g}(S[0\dots n])}}{g(n)}=\infty.$ (5.2)
Note that $\mathrm{KM}^{g}(x)=\mathrm{KM}^{g(n)^{2}}(x)+O(1)$}. Hence, if we assume that (5.1) is true then there is a $c\in\mathbb{R}$ and infinitely many $n$ with
$\mathrm{KM}^{g}(S[0\dots n])\leq\log\frac{1}{g(n)^{2}}+c.$
@@
$\frac{2^{-\mathrm{KM}^{g}(S[0\dots n])}}{g(n)^{2}}\geq 2^{-c}.$
Since $\lim_{n\to\infty}\frac{g(n)}{g(n)^{2}}=\infty$, (5.2) is true.
For the other direction we will prove the contrapositive. Suppose there is gauge $g$ such that $\mathrm{KM}^{g}(S[0\dots n])>\log\frac{1}{g(n)}$ for all sufficiently large $n$. Then we have
$\frac{2^{-\mathrm{KM}^{g}(S[0\dots n])}}{g(n)}\leq 1$
for all sufficiently large $n$ so (5.2) does not hold. ∎
Reimann and Slaman have defined [28] a class of sequences called never continuously random, denoted NCR, consisting of all sequences that are never effectively random with respect to any continuous probability measure. Here continuous means the same as atomless, but the outer measures for strong measure zero in (3.3) are more general than probability measures. The following is an overview of the definition of NCR, but see the paper for the details on their representations of measures.
@@
  • $X\in\mathbf{C}$ is in NCR if it its not $\mu$-random for every continuous probability measure $\mu$.
@@
First assume $S$ has algorithmic strong measure zero and consider any $Z$. Then for every unbounded increasing $h$ computable from $Z$ we have that $g(n)=2^{-h(n)}$} is a gauge also computable from $Z$. As $S$ has algorithmic strong measure zero, by Lemma 5.4 there is a constant $c$ and infinitely many n with
@@
Note that if there were only finitely many $n$ then an $h^{\prime}$ could be computed by changing finitely many values so that S is strongly $h^{\prime}$-complex. Thus, there are infinitely many $n$ with
@@ 6 Effective coverings and a Correspondence Principle
Let $\mathrm{\Delta}\in\mathrm{\Gamma}_{RCA}$, $A\subseteq\mathbb{N}$ and $X\subseteq\mathbf{C}$. A function $f^{A}:\mathbb{N}\to\{0,1\}^{*}$} is a $\mathrm{\Delta}$-AA covering of $X$ if $f\in\mathrm{\Delta}$ with oracle $A:\mathbb{N}\to\{0,1\}$ satisfies the following.
@@
We refer to $\{f^{A}(n)\mid n\in A\}$ of such an $f$ as a $\mathrm{\Delta}$-AA cover of $X$ and say that $X$ is strongly $\mathrm{\Delta}$-coverable if there is a functional $F\in\mathrm{\Delta}$ such that $F(A)$ is a $\mathrm{\Delta}$-AA covering of $X$ for all infinite $A\subseteq\mathbb{N}$.
@@
4: Run $F^{O}$ and for each $w$ with $|w|=f_{A}(n)$ where $F^{O}(w)\geq 1$ is found set c=c1c=c-1.
@@
Note that by construction the maximum number of strings $w$ with $|w|=f_{A}(n)$} and $F^{O}(w)>=1$ is at most $2^{n}$ in order for the supermartingale condition (3.1) to be satisfied. By construction, a prefix of each of these $w$ is output for every $n$. Moreover, for every $S\in X$, $F^{O}(S[0...n])$ can only increase on the lengths in $\mathrm{Range}(f_{A})$. Therefore $M^{A}$ will output a prefix of $S$ for every $S$ with $\limsup_{n\to\infty}F^{O}(S[0\dots n])>1$ and hence all $S\in X.$
@@
$f_{O}(n)=\min\{m\in\mathbb{N}\mid\prod_{i=0}^{m-1}O(w[0\dots i])\geq n\text{ for all }w\in\{0,1\}^{m}\}$
@@
Then $d_{i}$ is a $\Delta$-computable $O$-martingale. For each $w$ in the range of $M^{A_{i}}$ we have
@@
Let $\mathrm{\Delta}\in\mathrm{\Gamma}$. Then for every $X\subseteq\mathbf{C}$ there is an odds functional $F$ succeeding on $X$ if and only if there is an odds functional $F^{\prime}$ succeeding on $X$ where $F^{\prime}(O)$ is an odds martingale for all odds functions $O$. Moreover, there is a universal algorithmic martingale functional.
@@
  • $X\in\mathrm{\Sigma}_{2}^{0}$ if there is a computable function $f:\mathbb{N}\times\{0,1\}^{*}\to\{0,1\}$ where $T_{i}=\{w\mid f(i,w)=1\}$} are computable trees and $X=\bigcup_{i\in\mathbb{N}}X_{i}$} is the union of the $\mathrm{\Pi}_{1}^{0}$ sets corresponding to the $T_{i}$’s.
@@
Moreover 1 and 2 are equivalent for every union of $\mathrm{\Pi}_{1}^{0}$ sets.
@@
for every computable ordinal $\alpha$ there is a $x$ with $x\equiv_{\mathrm{T}}\emptyset^{(\alpha)}$ and $\{x\}$ having computable strong measure zero.